# Solvability of fractional semilinear heat equations with distributional forcing terms in Lorentz spaces

### v1

@@ 3.2 Contraction mapping argument
$\leqslant\int_{0}^{t}\left\|S(t-\tau)\mu\mathrel{}\middle|\mathrel{}B^{s-N/p}_{\infty,\infty}\right\|\left\|\psi\mathrel{}\middle|\mathrel{}B^{N/p-s}_{1,1}\right\|\,d\tau$
$\quad+\int_{0}^{t}\left\|S(t-\tau)\left[|u|^{\gamma-1}u(\cdot,\tau)\right]\mathrel{}\middle|\mathrel{}B^{s-N/p}_{\infty,\infty}\right\|\left\|\psi\mathrel{}\middle|\mathrel{}B^{N/p-s}_{1,1}\right\|\,d\tau$
$\leqslant\int_{0}^{t}\left\|S(t-\tau)\mu\mathrel{}\middle|\mathrel{}B^{s}_{(p,\infty),\infty}\right\|\,d\tau\left\|\psi\mathrel{}\middle|\mathrel{}B^{N/p-s}_{1,1}\right\|$
$\quad+\int_{0}^{t}\left\|S(t-\tau)\left[|u|^{\gamma-1}u(\cdot,\tau)\right]\mathrel{}\middle|\mathrel{}B^{\max\left\{0;s+\frac{(\gamma-1)N}{p}\right\}}_{(p/\gamma,\infty),\infty}\right\|\,d\tau\left\|\psi\mathrel{}\middle|\mathrel{}B^{N/p-s}_{1,1}\right\|$
$\leqslant C\left\|\psi\mathrel{}\middle|\mathrel{}B^{N/p-s}_{1,1}\right\|\bigg(\int_{0}^{t}1\,d\tau\left\|\mu\mathrel{}\middle|\mathrel{}B^{s}_{(p,\infty),\infty}\right\|$
$\qquad+\int_{0}^{t}(t-\tau)^{\min\left\{0;-\frac{s}{\theta}-\frac{(\gamma-1)N}{p\theta}\right\}}\left\||u|^{\gamma-1}u(\cdot,\tau)\mathrel{}\middle|\mathrel{}B^{0}_{(p/\gamma,\infty),\infty}\right\|\,d\tau\bigg)$
$\leqslant C\left\|\psi\mathrel{}\middle|\mathrel{}B^{N/p-s}_{1,1}\right\|\big(\left\|\mu\mathrel{}\middle|\mathrel{}B^{s}_{(p,\infty),\infty}\right\|+\left\|u\mathrel{}\middle|\mathrel{}X_{T}\right\|^{\gamma}\big)t^{\min\left\{1;1-\frac{s}{\theta}-\frac{(\gamma-1)N}{p\theta}\right\}}.$
Conditions $p\geqslant\frac{N(\gamma-1)}{\theta}$ and $s<0$ imply that the last term disappears in the limit $t\to 0$.
@@
Let $p>\gamma$ and $p\geqslant\frac{N(\gamma-1)}{\theta}$. Assume that a function $u=u(x,t)\in L^{\infty}(0,T;L^{p,\infty}_{\mathrm{ul}}(\mathbb{R}^{N}))$ satisfies the integral equation (1.2) for almost all $(x,t)\in\mathbb{R}^{N}\times\left]0,T\right[$, i.e., we have
@@
for almost all $(x,t)\in\mathbb{R}^{N}\times\left]0,T\right[$. The same calculation of the proof of Lemma 3.3 (estimate on $J[u]$) claims that $J[u](x,t)\in L^{\infty}(0,T;L^{p,\infty}_{\mathrm{ul}}(\mathbb{R}^{N}))$ and thus we deduce that
$I[\mu](x,t)=u(x,t)-J[u](x,t)$ $\in L^{\infty}(0,T;L^{p,\infty}_{\mathrm{ul}}(\mathbb{R}^{N})).$
Finally we apply Proposition 2.16 to derive $\left\|\mu\mathrel{}\middle|\mathrel{}B^{-\theta}_{(p,\infty),\infty}\right\|<\infty$. ∎
@@ Acknowledgements.
The author would like to thank Professor Ryo Takada for several comments. The author was supported in part by JST SPRING, Grant Number JPMJSP2108, and KAKENHI #24KJ0623.
Appendix A (2.13) for general $0<\theta\leqslant 2$
@@
$\left\|\mathcal{F}^{-1}\left[\Phi_{(0)}(\xi)C_{T}(\xi)\right]\mid L^{1}\right\|<\infty$ (A.1)
is valid for all $\theta\in\left]0,2\right]$, where
@@
$\mathcal{F}^{-1}\left[\Phi_{(0)}(\xi)C_{T}(\xi)\right]$
$=\frac{1}{(2\pi)^{N}}\int_{\mathbb{R}^{N}}e^{\sqrt{-1}x\xi}\rho\left(\frac{\xi}{\delta}\right)L^{\ast}(x,D)\left(\Phi_{(0)}(\xi)C_{T}(\xi)\right)\,d\xi$
@@
holds for all $k\geqslant 1$ and for all $\xi$. Note that we know $\left|\frac{d^{j}}{ds^{j}}\Psi\left(Ts\right)\right|\leqslant T^{j}C$ for bounded $s$ and for $j\in\mathbb{N}$. Thus we have the estimate
@@
$\left|\mathcal{F}^{-1}\left[\Phi_{(0)}(\xi)C_{T}(\xi)\right](x)\right|\leqslant\frac{CT^{-1}}{\left(1+|x|\right)^{N+\theta}}+\frac{CT^{-2}}{\left(1+|x|\right)^{N+1}}\leqslant\frac{CT^{-2}}{\left(1+|x|\right)^{N+\min\{1;\theta\}}}.$
Thus we conclude that (2.13) holds for general $0<\theta\leqslant 2$. ∎
@@
This attempt fails if we try to do so with $\Psi(T|\xi|^{\theta})\Phi_{j}(\xi)$ for all $j\in\mathbb{Z}$.

### v2

@@ 3.2 Contraction mapping argument
$\leqslant\int_{0}^{t}\left\|S(t-\tau)\mu\mathrel{}\middle|\mathrel{}B^{s-N/p,\infty}_{\infty}\right\|\left\|\psi\mathrel{}\middle|\mathrel{}B^{N/p-s,1}_{1}\right\|\,d\tau$
$\quad+\int_{0}^{t}\left\|S(t-\tau)\left[|u|^{\gamma-1}u(\cdot,\tau)\right]\mathrel{}\middle|\mathrel{}B^{s-N/p,\infty}_{\infty}\right\|\left\|\psi\mathrel{}\middle|\mathrel{}B^{N/p-s,1}_{1}\right\|\,d\tau$
$\leqslant\left\|\psi\mathrel{}\middle|\mathrel{}B^{N/p-s,1}_{1}\right\|\bigg(\int_{0}^{t}\left\|S(t-\tau)\mu\mathrel{}\middle|\mathrel{}B^{s,\infty}_{p,q,\infty}\right\|\,d\tau$
$\qquad+C\int_{0}^{t}\left\|S(t-\tau)\left[|u|^{\gamma-1}u(\cdot,\tau)\right]\mathrel{}\middle|\mathrel{}B^{\max\left\{0,s+\frac{(\gamma-1)N}{p}\right\},\infty}_{p/\gamma,q/\gamma,\infty}\right\|\,d\tau\bigg)$
$\leqslant C\left\|\psi\mathrel{}\middle|\mathrel{}B^{N/p-s,1}_{1}\right\|\bigg(\int_{0}^{t}1\,d\tau\left\|\mu\mathrel{}\middle|\mathrel{}B^{s,\infty}_{p,q,\infty}\right\|$
$\qquad+\int_{0}^{t}(t-\tau)^{\min\left\{0,-\frac{s}{\theta}-\frac{(\gamma-1)N}{p\theta}\right\}}\left\||u|^{\gamma-1}u(\cdot,\tau)\mathrel{}\middle|\mathrel{}B^{0,\infty}_{p/\gamma,q/\gamma,\infty}\right\|\,d\tau\bigg)$
$\leqslant C\left\|\psi\mathrel{}\middle|\mathrel{}B^{N/p-s,1}_{1}\right\|\left(\left\|\mu\mathrel{}\middle|\mathrel{}B^{s,\infty}_{p,q,\infty}\right\|+\left\|u\mathrel{}\middle|\mathrel{}X_{T}\right\|^{\gamma}\right)t^{\min\left\{1,1-\frac{s}{\theta}-\frac{(\gamma-1)N}{p\theta}\right\}}.$
Conditions $p\geqslant{N(\gamma-1)}/{\theta}$ and $s<0$ imply that the last term disappears in the limit $t\to 0$.
@@
Let $p\geqslant q>\gamma$ and $p\geqslant{N(\gamma-1)}/{\theta}$. Assume that a function $u=u(x,t)\in L^{\infty}(0,T;M^{p}_{q,\infty}(\mathbb{R}^{N}))$ satisfies the integral equation (1.1) for almost all $(x,t)\in\mathbb{R}^{N}\times\left]0,T\right[$, i.e., we have
@@
for almost all $(x,t)\in\mathbb{R}^{N}\times\left]0,T\right[$. The same calculation of the proof of Lemma 3.3 (estimate on $J[u]$) leads $J[u](x,t)\in L^{\infty}(0,T;M^{p}_{q,\infty}(\mathbb{R}^{N}))$, hence the inclusion
$I[\mu](x,t)=u(x,t)-J[u](x,t)$ $\in L^{\infty}(0,T;M^{p}_{q,\infty}(\mathbb{R}^{N})).$
Finally we apply Proposition 2.18 to derive $\left\|\mu\mathrel{}\middle|\mathrel{}B^{-\theta,\infty}_{p,q,\infty}\right\|<\infty$. ∎
@@ Acknowledgements.
The author was supported in part by JST SPRING, Grant Number JPMJSP2108, and KAKENHI #24KJ0623.
Appendix A Estimate (2.13) for general $\theta>0$
@@
$\left\|\mathcal{F}^{-1}\left[\Phi_{(0)}(\xi)C_{T}(\xi)\right]\mathrel{}\middle|\mathrel{}L^{1}\right\|<\infty$ (A.1)
is valid for all $\theta>0$, where
@@
$\mathcal{F}^{-1}\left[\Phi_{(0)}(\xi)C_{T}(\xi)\right]$ $=\frac{1}{(2\pi)^{N}}\int_{\mathbb{R}^{N}}e^{\sqrt{-1}x\xi}\rho\left(\frac{\xi}{\delta}\right)L^{\ast}(x,D)\left(\Phi_{(0)}(\xi)C_{T}(\xi)\right)\,d\xi$
@@
holds for all $k\geqslant 1$ and for all $\xi$. Note that $\left|\frac{d^{j}}{ds^{j}}\Psi\left(Ts\right)\right|\leqslant T^{j}C$ for bounded $s$ and for $j\in\mathbb{N}$. Thus we have the estimate
@@
$\left|\mathcal{F}^{-1}\left[\Phi_{(0)}(\xi)C_{T}(\xi)\right](x)\right|\leqslant\frac{CT^{-1}}{\left(1+|x|\right)^{N+\theta}}+\frac{CT^{-2}}{\left(1+|x|\right)^{N+1}}\leqslant\frac{CT^{-2}}{\left(1+|x|\right)^{N+\min\{1,\theta\}}}.$
Thus we conclude that (2.13) holds for general $\theta>0$. ∎
@@
This attempt fails if we do with $\left\{\Psi(T|\xi|^{\theta})\Phi_{j}(\xi)\right\}_{j\in\mathbb{Z}}$.
Appendix B On the local Morrey-type spaces
In this section, we make some comments on the local Morrey-type spaces.
Proposition B.1.
Let $1\leqslant q_{0}<q_{1}<p<\infty$. Then the continuous embeddings $M^{p}_{q_{0}}\supset M^{p}_{q_{1}}\supset L^{p,\infty}_{\mathrm{ul}}$ hold and these inclusions are proper, i.e., there exists some functions $f$ such that $f\in M^{p}_{q_{0}}$ but $f\not\in M^{p}_{q_{1}}$. Here $L^{p,\infty}_{\mathrm{ul}}$ is defined as the set of measurable function $f$ on $\mathbb{R}^{N}$ satisfying
$\left\|f\mathrel{}\middle|\mathrel{}L^{p,\infty}_{\mathrm{ul}}\right\|:=\sup_{z\in\mathbb{R}^{N}}\left\|f\chi_{B(z,1)}\mathrel{}\middle|\mathrel{}L^{p,\infty}\right\|<\infty.$
Proof.
The embedding $M^{p}_{q_{0}}\supset M^{p}_{q_{1}}$ comes from the Hölder inequality. We can prove the latter $M^{p}_{q_{1}}\supset L^{p,\infty}_{\mathrm{ul}}$ in the same manner as Lemma 1.7 in [20]. Thus it is enough to give an example $f\in M^{p}_{q_{0}}\setminus M^{p}_{q_{1}}$. We modify the proof of Theorem 1.3 (i) in [10]. For any $c>d>0$, we define the annulus $A^{c}_{d}:=B(0,c)\setminus B(0,d)$, and for each $k\in\mathbb{Z}_{\geqslant 1}$ and any $\delta>0$, we define the following quantities:
$Q_{0}(k):=\int_{A^{1/k}_{1/(k+1)}}|x|^{-\frac{Nq_{0}}{p}}\,dx,\quad Q_{1}(k):=\int_{A^{1/k}_{1/(k+1)}}|x|^{-\frac{Nq_{1}}{p}-\delta}\,dx,$
$a(k):=\max\left\{\left(\frac{Q_{1}(k)}{Q_{0}(k)}\right)^{\frac{1}{q_{1}-q_{0}}},\left(\frac{Q_{1}(k)}{\left|A^{1/k}_{1/(k+1)}\right|}\right)^{1/q_{1}}\right\}.$
The following two inequalities immediately follow from the definition of $a(k)$.
$\frac{Q_{1}(k)}{a(k)^{q_{1}}}\leqslant\frac{Q_{0}(k)}{a(k)^{q_{0}}},\quad\left(0<\right)\,\frac{Q_{1}(k)}{a(k)^{q_{1}}}\leqslant\left|A^{1/k}_{1/(k+1)}\right|.$ (B.1)
Since the function $\varepsilon\mapsto r_{k}(\varepsilon):=\left|A^{1/(k+1)+\varepsilon}_{1/(k+1)}\right|$ is increasing and continuous, and satisfies $r_{k}(0)=0$ and $r_{k}\left(1/k-1/(k+1)\right)=\left|A^{1/k}_{1/(k+1)}\right|$, we can define the quantity $\varepsilon(k)\in\left]0,1/k-1/(k+1)\right]$ by the formula $\left|A^{1/(k+1)+\varepsilon(k)}_{1/(k+1)}\right|=\dfrac{Q_{1}(k)}{a(k)^{q_{1}}}$. Now we define a function $f:\mathbb{R}^{N}\to\mathbb{R}$ by
$f(x):=\sum_{k=1}^{\infty}a(k)\chi_{A^{1/(k+1)+\varepsilon(k)}_{1/(k+1)}}(x).$
We check that this $f$ satisfies $f\in M^{p}_{q_{0}}\setminus M^{p}_{q_{1}}$. Let $m=m(R)$ denote the unique integer which satisfies $1/(m+1)<R\leqslant 1/m$, for each $R\in\left]0,1\right]$. Then we have
$\int_{B(0,R)}|f(x)|^{q_{0}}\,dx$ $=\sum_{k=m(R)}^{\infty}a(k)^{q_{0}}\left|A_{1/(k+1)}^{1/(k+1)+\varepsilon(k)}\right|=\sum_{k=m(R)}^{\infty}a(k)^{q_{0}}\frac{Q_{1}(k)}{a(k)^{q_{1}}}$
(B.1) $\leqslant\sum_{k=m(R)}^{\infty}a(k)^{q_{0}}\frac{Q_{0}(k)}{a(k)^{q_{0}}}\leqslant\int_{B(0,2R)}|x|^{-\frac{Nq_{0}}{p}}\,dx=CR^{-\frac{Nq_{0}}{p}+N}$
and thus for arbitrary $R\in\left]0,1\right]$, the estimate
$R^{\frac{N}{p}-\frac{N}{q_{0}}}\left\|f\chi_{B(0,R)}\mathrel{}\middle|\mathrel{}L^{q_{0}}\right\|\leqslant CR^{\frac{N}{p}-\frac{N}{q_{0}}}R^{-\frac{N}{p}+\frac{N}{q_{0}}}=C<+\infty$
holds. Similarly, for $\delta\in\left]0,N\left(1-\frac{q_{1}}{p}\right)\right[$, we see that
$\int_{B(0,R)}|f(x)|^{q_{1}}\,dx=\sum_{k=m(R)}^{\infty}a(k)^{q_{1}}\frac{Q_{1}(k)}{a(k)^{q_{1}}}\geqslant\int_{B(0,R)}|x|^{-\frac{Nq_{1}}{p}-\delta}\,dx=CR^{-\frac{Nq_{1}}{p}-\delta+N}.$
Therefore, sending $R\to 0$, we deduce
$R^{\frac{N}{p}-\frac{N}{q_{1}}}\left\|f\chi_{B(0,R)}\mathrel{}\middle|\mathrel{}L^{q_{1}}\right\|\geqslant CR^{\frac{N}{p}-\frac{N}{q_{1}}}R^{-\frac{N}{p}-\frac{\delta}{q_{1}}+\frac{N}{q_{1}}}=CR^{-\frac{\delta}{q_{1}}}\rightarrow+\infty,$
which means $f\not\in M^{p}_{q_{1}}$. ∎
The function $f$ constructed above does not satisfy that $f(x)\leqslant a|x|^{-N/p}$ for $x\in B(0,\rho)$, for any $a>0$ and $\rho>0$. Finally, we remark that considering “Morrey–Morrey” spaces is not meaningful.
Proposition B.2.
Let $1\leqslant r\leqslant q\leqslant p<\infty$. Then we have $M^{p}_{r}=M^{p}_{M^{q}_{r}}$. Here,
$M^{p}_{M^{q}_{r}}:=\left\{f\in L^{1}_{\mathrm{loc}}(\mathbb{R}^{N})\mathrel{}\middle|\mathrel{}\left\|f\mathrel{}\middle|\mathrel{}M^{p}_{M^{q}_{r}}\right\|:=\sup_{\begin{subarray}{c}0<R\leqslant 1\\ z\in\mathbb{R}^{N}\end{subarray}}R^{\frac{N}{p}-\frac{N}{q}}\left\|f\chi_{B(z,R)}\mathrel{}\middle|\mathrel{}M^{q}_{r}\right\|<\infty\right\}.$
Proof.
Fix $R\in\left]0,1\right]$ and $z\in\mathbb{R}^{N}$. For all $f\in M^{p}_{M^{q}_{r}}$, we have the inequality
$R^{\frac{N}{p}-\frac{N}{r}}\left\|f\chi_{B(z,R)}\mathrel{}\middle|\mathrel{}L^{r}\right\|$ $=R^{\frac{N}{p}-\frac{N}{q}}R^{\frac{N}{q}-\frac{N}{r}}\left\|f\chi_{B(z,R)}\chi_{B(z,R)}\mathrel{}\middle|\mathrel{}L^{r}\right\|$
$\leqslant R^{\frac{N}{p}-\frac{N}{q}}\left\|f\chi_{B(z,R)}\mathrel{}\middle|\mathrel{}M^{q}_{r}\right\|\leqslant\left\|f\mathrel{}\middle|\mathrel{}M^{p}_{M^{q}_{r}}\right\|,$
and this implies $M^{p}_{r}\supset M^{p}_{M^{q}_{r}}$. Next we take any $f\in M^{p}_{r}$ and consider the function
$Q:\left]0,1\right]^{2}\times\mathbb{R}^{N+N}$ $\rightarrow\mathbb{R},$
$\left(R,\tilde{R},z,w\right)$ $\mapsto Q\left[R,\tilde{R},z,w\right]:=R^{\frac{N}{p}-\frac{N}{q}}\tilde{R}^{\frac{N}{q}-\frac{N}{r}}\left\|f\chi_{B(z,R)}\chi_{B(w,\tilde{R})}\mathrel{}\middle|\mathrel{}L^{r}\right\|.$
Assume $R\geqslant\tilde{R}$. Then $p\geqslant q$ tells us $\left(R/\tilde{R}\right)^{\frac{N}{p}-\frac{N}{q}}\leqslant 1$ and thus we get
$Q\left[R,\tilde{R},z,w\right]\leqslant\left(\frac{R}{\tilde{R}}\right)^{\frac{N}{p}-\frac{N}{q}}\tilde{R}^{\frac{N}{p}-\frac{N}{r}}\left\|f\chi_{B(w,\tilde{R})}\mathrel{}\middle|\mathrel{}L^{r}\right\|\leqslant\left\|f\mathrel{}\middle|\mathrel{}M^{p}_{r}\right\|.$
Conversely assume $R\leqslant\tilde{R}$. Then $r\leqslant q$ tells us $\left(R/\tilde{R}\right)^{\frac{N}{r}-\frac{N}{q}}\leqslant 1$ and thus we have
$Q\left[R,\tilde{R},z,w\right]\leqslant\left(\frac{R}{\tilde{R}}\right)^{\frac{N}{r}-\frac{N}{q}}R^{\frac{N}{p}-\frac{N}{r}}\left\|f\chi_{B(z,R)}\mathrel{}\middle|\mathrel{}L^{r}\right\|\leqslant\left\|f\mathrel{}\middle|\mathrel{}M^{p}_{r}\right\|.$
Combining the above two yields $Q\leqslant\left\|f\mathrel{}\middle|\mathrel{}M^{p}_{r}\right\|$ and this finishes the proof. ∎